### v1

# Keeping Score: A Quantitative Analysis of How the CHI Community Appreciates Its Milestones

Abstract.
The ACM CHI Conference has a tradition of citing its intellectual heritage. At the same time, we know CHI is highly diverse and evolving. In this highly dynamic context, it is not clear how the CHI community continues to appreciate its milestones (within and outside of CHI). We present an investigation into how the community’s citations to milestones have evolved over 43 years of CHI Proceedings (1981–2024). Forgetting curves plotted for each year suggest that milestones are slowly fading from the CHI community’s collective memory. However, the picture is more nuanced when we trace citations to the top-cited milestones over time. We identify three distinct types of milestones cited at CHI, a typology of milestone contributions, and define the Milestone Coefficient as a metric to assess the impact of milestone papers on a continuous scale. Further, we provide empirical evidence of a Matthew effect at CHI. We discuss the broader ramifications for the CHI community and the field of HCI.
quantitative analysis, milestones, CHI, citations, forgetting curves, bibliometric analysis, Matthew effect, meta-science, meta-HCI.
journalyear: 2024
Figure 1. Forgetting curves (Reisz et al., 2022) for the Proceedings of the ACM Conference on Human Factors in Computing Systems (CHI) from 1981 to 2024. A forgetting curve plots the number of citations CHI papers make to prior papers published during the years plotted on the x-axis. The primary y-axis represents the citations in these publications over time, adjusted for overall growth in the number of publications in the proceedings, and scaled logarithmically for comparison. Each line represents the forgetting curve for papers published in the year following the right-most point of the line. For instance, the curve at the bottom is the forgetting curve for the 2024 CHI proceedings. The peaks in the forgetting curves visually indicate the presence of “milestones”, when read along a vertical line from top to bottom. In other words, several proceedings contain particularly many references to milestones published in that year. For forgotten milestones, this visible peak is lost over time. Here, ‘forgetting’ refers to a decrease in citations over time. We can see how the proceedings of the last few years are distributing their citations more uniformly, with less spikes and dips visible in the curves. This is indicative of the community’s divided attention outside what was in earlier years considered as its key milestones. Independent of the forgetting curves, the secondary y-axis depicts the number of citations from all CHI papers ever published to the top-30 milestones most cited at CHI (see Table 1) per their publication years (x-axis). For instance, Braun and Clarke’s milestone in the year 2006 (Braun and Clarke, 2006) received many citations and the adoption of this milestone in the CHI community is visible as a growing peak in the forgetting curves. The fact that there were no proceedings in 1984 contributes to fluctuations in the older curves.
[…]
Forgetting curves for citations in the Proceedings of the ACM CHI Conference on Human Factors in Computing Systems from 1981 to 2024.
1. Introduction
Human-Computer Interaction (HCI) is highly interdisciplinary and characterized by constant evolution and changes (Fuchsberger et al., 2021; Bødker, 2015). Recent changes include, to name but a few, an overall growth in the number of scientific papers published each year (Fortunato et al., 2018; Bornmann et al., 2021), and an acceleration of pace in science with mounting career and publication pressures (Edwards and Roy, 2017; Oppenlaender, 2024). In addition, sudden technological advances can shift focus and topics in HCI research. One such technological advance is the recent progress in large auto-regressive language models, such as ChatGPT (Bubeck et al., 2023; Vaswani et al., 2017). The increasing integration of language models into HCI research reflects broader trends in methodological acceleration and automation (Thomas et al., 2024). While novel AI tools can streamline certain tasks, over-reliance on such tools and their misuse has led to concerns about prompt-hacking (Morris, 2024) and shallow inquiries that prioritize speed over depth (Thomas et al., 2024; Kobak et al., 2024; Gray, 2024; Liang et al., 2024; Academ-AI, 2024; Arawjo, 2024). As HCI evolves, it is critical to reflect on the types of contributions that drive long-lasting impact, particularly amid shifting research norms and topics. In light of this dynamic context, it is important to explore what we value in HCI and understand how the CHI community interacts with its intellectual heritage.
The CHI community has a history of citing and valuing its intellectual heritage. For instance, the “Guide to a Successful Paper” at CHI ’16 expressively called authors to “make sure to cite prior work […] in the relevant area” and works “that have had a major influence on your own work” (ACM SIGCHI, 2016). Letting others build on one’s work has been quoted to be “the entire purpose of a CHI Paper” and “lack of references” was identified as “a frequent cause for complaint – and low rating – by reviewers” (ACM SIGCHI, 2016). These sentiments help steer CHI, highlighting the community’s commitment to acknowledging relevant prior work. The emphasis on recognizing intellectual heritage highlights the importance of understanding which works have shaped the field of HCI. However, beyond subjective appreciation, a more systematic approach is needed to quantify the influence of these impactful works.
Citation analysis, while not the only one, is the de facto means for analyzing the impact of past contributions. In this work, we explore quantitatively how the CHI community values its past milestones: articles that have attained exceptional attention from the community. These articles are important to the community, and thus to the discipline itself. For instance, the System Usability Scale (SUS) (Brooke, 1996) and the NASA-TLX (Hart and Staveland, 1988) are two widely adopted measurement scales used in HCI studies. Given the interdisciplinarity of the field, these milestones do not always originate within the community. Thematic Analysis guidelines (Braun and Clarke, 2006) are a great example of an influential contribution that did not originate from CHI or even the broader HCI community. We explore how the CHI community pays attention to its milestones over time, exploring particularly the following research questions:
  • RQ1:
    What are the milestones most cited in the CHI community?
  • RQ2:
    What are the contributions of these milestones?
  • RQ3:
    How has the CHI community’s attention to its milestones evolved over time?
  • RQ4:
    Does the CHI community show a preference for its milestone authors?
We present the results of a quantitative analysis of 11,542 articles published in the Proceedings of the ACM CHI Conference on Human Factors in Computing Systems (1981–2024, with the exception of 1984). Our citation curve analysis examines the CHI community’s current attention span and the community’s selective memory. The majority of milestone citations point 5–15 years back in time, suggesting a “sweet spot” of relevance for papers published at CHI. As expected, we also find that the community milestones make diverse types of contributions. Yet, a comparison with existing well-known research contribution taxonomies (Wobbrock and Kientz, 2016; ACM SIGCHI, 2024), reveals that popular CHI contribution types (such as prototypes and systems), but also other contribution types (such as surveys, datasets, and replications), are absent in the CHI community’s milestones. This highlights how some highly popular contribution types—such as prototypes and systems—all play a role but seldom leave a long-lasting historical mark. We present the 30 milestones most cited by the CHI community in this paper, and we make a dataset of the 300 most cited milestones available in the spirit of Open Science. [^1] Link to Zenodo will be provided after anonymous peer review. We further present a case study on CHI’s “super milestone”, Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006), that continues to enjoy extraordinary popularity at CHI. Finally, we present the Milestone Coefficient as a metric for assessing research impact of milestones. This metric offers a more nuanced understanding of what constitutes a milestone beyond mere citation count. We discuss our findings in light of related literature and provide our reflections on how the community remembers, forgets, and curates past knowledge.
Our research contributes to meta research on HCI literature, particularly focusing on the CHI community, as a not fully representative but important community in the field of HCI.
2. Related Work
While the HCI field is buzzing with activity, as demonstrated by the well-visited and growing CHI Conference, surprisingly little research and self-reflection on HCI research is being published at CHI. For example, the HCI field’s top conference currently has no subcommittee for meta-scientific investigations. This contrasts with the field’s critical and self-reflective tradition (Sengers et al., 2005; Schön, 1992). Such self-reflective and meta-scientific contributions are critically important to advance the HCI field as a whole, and in the following we present prior meta-research on HCI and CHI.
2.1. Bibliometric Analyses at CHI
Pohl and Mottelson (Pohl and Mottelson, 2019) conducted a quantitative analysis of 6,578 CHI papers, investigating how authors write their papers and how factors such as readability and name dropping influence citation counts. Bartneck and Hu (Bartneck and Hu, 2009) presented a scientometric analysis of CHI proceedings, focusing on organizations and countries that contribute to CHI. Their work highlighted the difficulty of judging quality in the context of best paper awards, finding a mismatch between awarded papers and citation counts. Another bibliometric analysis was presented by Lee et al.(Lee et al., 2019). Their citation network analysis and cluster analysis identified emerging research themes in HCI. Research themes were also the topic of Oppenlaender and Hämäläinen’s work on research challenges in HCI (Oppenlaender and Hämäläinen, 2023). The authors analyzed CHI papers with the use of language models and mapped out the diverse research landscape of CHI. A more traditional approach to analyzing topics in CHI was taken by Liu et al. who presented a co-word analysis of the evolution of the HCI field (Liu et al., 2014). Their thematic analysis of keywords in CHI papers published between 1994 and 2013 found a fragmentation of research approaches within HCI. Finally, Kaye (Kaye, 2009) investigated repeat authorship at CHI. Among other findings, Kaye found the mean number of authors in CHI papers is increasing, and many CHI authors are repeat authors.
It can be noted that the above approaches did neither consider the time dimension of citation networks nor preferential attachment (Barabási and Albert, 1999; Merton, 1968), a self-reinforcing mechanism where successful papers become more successful over time. These factors concern the dynamic nature of scholarly publishing, where older papers have more time to accrue an interest in the community. Therefore, a citation analysis must always consider the growth in the number of publications (Reisz et al., 2022). These dynamics and mechanisms play a key role in our analysis of the CHI community’s milestones.
2.2. Milestones
Human-Computer Interaction is highly interdisciplinary, borrowing methods and approaches from diverse fields such as Design, Psychology, and Computer Science. Perhaps it is for this interdisciplinarity and HCI being an “inter-discipline” (Blackwell, 2015) that milestones have, so far, not been a topic in the HCI literature. Milestones are highly cited papers that move a field forward. In many cases, these papers invented new methods, frameworks, or entire sub-fields in HCI that have inspired hundreds of research papers. We feel strongly that milestone is the appropriate name for such papers. Moreover, the term milestone is also used in other fields in a similar fashion. Being highly cited over years, milestone papers show resilience against aging. In this context, aging of a publication refers to the preference of authors to cite more recent works over older works (Reisz et al., 2022; Wang et al., 2013; Yin and Wang, 2017).
However, there is no standard definition of what constitutes a milestone paper. Each scientific field moves at a different pace, and even within a given scientific field or discipline, there is no clear definition of what constitutes a milestone. This is evident in the different approaches taken by authors in prior research to identify milestone papers. Redner used the total number of citations over a paper’s lifespan as criteria, focusing on papers with more than 1,000 citations (Redner, 2005). On the other hand, Newman examined the 50 best-performing papers (Newman, 2014), and Reisz et al. focused on the top-30 most-cited papers (Reisz et al., 2022). Ke et al. took an entirely different approach, focusing on wake-up intensity of would-be milestone papers (“sleeping beauties”) (Ke et al., 2015). Common to these approaches is that a milestone paper is a work that receives an exceptional number of citations, compared to other papers in a given research field. In the following section, we review related work on a research community’s attachment to its milestones.
2.3. Citation Memory
A research community’s collective “memory” is expressed through its citations to past works. Preferential attachment (Barabási and Albert, 1999) is the phenomenon that highly cited papers are more visible and, thus, more likely to be cited than less-cited papers (Barabási and Albert, 1999; Wang et al., 2013). The preferential attachment of a research community to its high-performing papers (and authors) is also referred to as accumulated advantage or Matthew effect in science (Merton, 1968).
On the other hand, past works may fall out of favor due to factors such as technological progress and shifts in research topics over time. In these cases, the research community’s collective memory may be fading. In (1960), Burton and Kebler coined the term half-life of academic literature to describe this phenomenon (Burton and Kebler, 1960). A related concept is citation amnesia, which refers to how far back in time researchers tend to cite (Singh et al., 2023).
It can be concluded that the number of citations a paper receives depends both on the paper’s accrued citations and its age (Wang et al., 2013; Yin and Wang, 2017). Ultimately, the absolute number of citations is not the only imaginable indicator for a paper’s impact (Newman, 2014), and citation analysis should always consider the dynamics and exponential growth of the number of publications (Reisz et al., 2022; Wang et al., 2013; Yin and Wang, 2017). In the following section, we describe how forgetting curves are a means of analysis that considers this growth in publications.
2.4. Forgetting Curves
A citation curve plots the number of citations to past papers from papers published in a given year. For instance, if a paper published in the year 2024 cites a paper published in 1950, this counts as one citation to the 1950 paper. Citation curves provide a visual indication of the attention span of authors. On the one hand, a flat citation curve indicates a long community memory where many old papers are being cited in a given proceedings year. On the other hand, a citation curve with a steep and right-leaning peak indicates a short community memory where old papers are quickly being forgotten. However, one factor that comes into play is the number of articles published each year. This number has been increasing in recent years, thereby introducing a bias to the insights that we can derive from citation curves.
Forgetting curves (Reisz et al., 2022) address this problem. A forgetting curve is a growth-adjusted citation curve. For all references in a given proceedings year, the forgetting curve looks at what prior years have been cited and counts citations to these years (i.e., the citation curve). The forgetting curve is calculated as follows:
(1) $\bar{c}_{t_{i}t_{j}}=\frac{N_{t_{j}}}{N}\sum_{t_{j}}c_{t_{i}t_{j}}$
where $N_{t_{j}}$ is the number of papers published in year $t_{j}$, $N$ is the total number of papers published in all years, and $c_{t_{i}t_{j}}$ is the citation curve (Reisz et al., 2022). The sum $\sum_{t_{j}}c_{t_{i}t_{j}}$ is the total number of citations from papers published in year $i$ to papers published in year $j$. To mitigate bias due to growth in publications over the years, the forgetting curve is weighted with the relative number of publications. This weighting factor in formula (1) together with the citation counts can be interpreted as an expectancy value. Valleys in the forgetting curve will form when citations to a past year under-perform compared to the expected value, and peaks will indicate above average performance. See the work by Reisz et al. (Reisz et al., 2022) for a more detailed explanation of citation curves and forgetting curves.
Past papers that receive many citations from papers published in a given year will form visible peaks in the forgetting curve of this year. Plotting forgetting curves over several years, therefore, allows us to visually inspect how citations to milestone papers have evolved over time. To this end, the forgetting curve is read along a vertical line for a given year. If peaks form along this vertical line when read from top to bottom, milestone papers are being discovered and cited over time. On the other hand, if peaks vanish between the curves when read from top to bottom, it is an indication that milestone papers published in a given year are being “forgotten” in later years. Forgotten, here, refers to a decline in citations over time. Of course, the paper may still continue to receive citations (as most papers do over time). But in relative terms, it is being forgotten. In this work, we apply forgetting curves to the ACM CHI proceedings.
3. Method
3.1. Data Collection
The data for this analysis was collected from the ACM Digital Library (ACM–DL). For each research article in the CHI proceedings from 1981–2024 (with exception of 1984), we downloaded the article’s list of references, as listed on the article’s page in the ACM–DL. We exclude keynotes, panel sessions, and abstract only articles. All information collected for this study is publicly accessible on the ACM–DL. The resulting dataset consists of 11,542 articles and a total of 629,120 references. From this set of articles and references, we identify the milestone papers cited by the CHI community, as described in the following section.
3.2. Milestones
The following sections describe how we identify milestones, their types and contributions, and the milestone authors.
3.2.1. Milestone definition and identification
As mentioned in Section 2.2, there is no standard method for identifying milestone papers and different approaches have been used in prior work. In our work, the papers with the highest total number of citations from CHI papers (over all CHI proceedings) are considered the milestones of the CHI community. More specifically, we adopt the approach by Reisz et al. (Reisz et al., 2022) and define the top 30 most cited publications as milestones in our initial analysis. As in Reisz et al., we acknowledge this choice is arbitrary. However, to solidify our analysis we later extend our analysis to the top-100, top-300, and top-3000 papers cited in the CHI community, finding that the results are consistent for larger groups of milestones. We further define “super milestones” as high performing milestones which have received more than 0.1% (1 in 1000) of all citations in at least one proceedings year. We derived this threshold empirically from visual inspection of citations over time (cf. Figure 2). While this threshold is also arbitrary, it serves as a meaningful decision boundary for discussing different types of milestones in our analysis.
Note that in our analysis and throughout this whole article, milestones are not limited to being published at CHI, but can also appear in other publication venues. A milestone can be a conference paper, journal article, or whole book. For simplicity, we refer to them as ‘milestones’ or ‘milestone papers’ in the remainder of this work. Further note that the citation counts presented in our analysis (cf. Table 1) differ significantly from those that include all citations to a milestone, such as citation counts reported by Google Scholar. Our focus in this work rests exclusively on the CHI community and how it cites its milestones.
To identify the milestones, we parse the reference texts and extract the publication year and the title with an iteratively developed set of over 100 regular expressions and logical checks. We apply basic normalization to the paper’s title, including lowercasing as well as removal of spaces, line breaks, and special characters. By concatenating the year and normalized title into a unique ID, we can robustly identify papers by this ID (though, see Section 5.5 for limitations of this approach). To answer RQ1, we can then analyse the citations from CHI papers to the uniquely identified papers.
3.2.2. Milestone types and contributions
To answer RQ2, we identify different types of milestones and analyze their contributions. To identify types of milestones, we performed a cluster analysis, based on the data from the citation time series of the top-100 milestones most cited by CHI papers (over all proceedings years). We triangulated this problem with different clustering approaches, including Ward’s method (Jr., 1963) and dynamic time warping with k-means (Petitjean et al., 2011). The latter is a clustering technique that aligns and groups time series data with varying lengths by measuring similarity through optimal time alignment, allowing patterns to be identified despite temporal distortions. This technique was used for clustering milestones in Table 1 and Figure 2. We explored the optimal number of clusters with the elbow criteria (Thorndike, 1953) and other methods, as listed in (Schubert, 2023). However, analysis of dendrogram plots proved to provide the most reasonable and meaningful results, pointing to a three-cluster solution.
To identify the milestone contributions, we inductively coded the key contributions of the top-30 milestones most cited in the CHI community. To this end, the first author read the milestones’ title and abstract and iteratively coded the milestones. The codes were discussed in two meetings with the co-authors to reach consensus. This approach ensured consistency and alignment in the interpretation of the data. One code was assigned to each milestone paper. We acknowledge that a milestone may make more than one contribution. However, our coding aimed to capture the key contribution of the community’s top-performing milestones, answering the question of “what is the milestone most known and cited for?” The identified set of contributions was then mapped to existing contribution types by Wobbrock and Kientz (Wobbrock and Kientz, 2016) and ACM SIGCHI (ACM SIGCHI, 2024).
3.2.3. Author identification
We parse author names from the ACM–DL’s citation text and the scraped CHI paper references with a combination of regular expressions and logical checks to identify the author names in different reference notation styles (e.g., “Firstname Lastname. 2023.” or “Lastname, Initials (2023a)“). Typically, these notation styles are not mixed within papers and individual references. This allows us to relatively reliably detect the pattern of author names and extract the author names from the CHI papers’ references with regular expressions.
For uniquely identifying authors, we follow Kaye’s approach (Kaye, 2009) and base our analysis on only the author’s last name and the first initial. While this may lead to author names being conflated, abbreviating first names is necessary because it constitutes the lowest common denominator in the different references formats found in CHI papers. We discuss further limitations to this approach in Section 5.5.
3.3. Forgetting Curve Plots
We plot forgetting curves for the available CHI Proceedings years. The construction of forgetting curves is straight-forward from the scraped data, using the formula given in Section 2.4. The forgetting curves include citations to papers published at CHI and other venues. Therefore, the curves reach far back in time, but for the purpose of analyzing CHI, we limit the plot to our time frame of analysis (1981–2024). On the secondary axis (in Figure 1), we plot the total number of citations to the 30 most cited milestones in their respective publication year as vertical bars. The bars help in visually identifying some of the milestone peaks in the forgetting curves.
3.4. Milestone Coefficient
We define the milestone coefficient as a measure for the growing or diminishing impact of a paper relative to other papers published within a research community. The aim of this metric is to classify papers on a continuous open-ended “milestone spectrum.” We design the metric to amplify differences between milestones that were popular once, but are being forgotten, and milestones that continue to grow in citations. Our aim with this metric is to provide a tool for analyzing scientific papers cited by a research community. The metric goes beyond a discrete typology of milestones and instead uses a continuous spectrum to measure the success of a paper in a research community. To this end, we first define the trend of the citation series $c$ of a paper:
$\text{SIGN}=\begin{cases}1,&\text{if }m\geq 0\\ -1,&\text{if }m<0\end{cases}\text{with }m=\begin{cases}\text{slope of the linear regression of }c\text{ on }x,&\text{if }n>1\\ 0,&\text{if }n=1\end{cases}$
where $c$ is the citation series
$c=\{c_{0},c_{1},\dots,c_{n-1}\}$
and each $c_{i}$ is the citation count for year $x_{i}$ and $n$ is the number of years in the series. A milestone that is fading from the collective memory of the research community will have a negative SIGN value, indicating a gradual decline in citations to this milestone. One the other hand, a milestone that is growing in citations has a positive SIGN value. If there is only one citation value in the series ($n=1$), we define the slope $m=0$ and $SIGN=1$. A paper’s Total Citation Span (TCS) is a measure of the spread of citation counts over time. It is calculated as the difference between the highest and lowest citation counts observed for the paper:
$\text{TCS}=\max(c)-\min(c)$
where $c$ is the series of citations to the paper over time. TCS reflects the range in how frequently a paper has been cited, indicating the extent to which its citation count fluctuates. A higher TCS suggests greater variability in citation counts, while a lower TCS indicates a more consistent citation frequency. We adjust the TCS by the paper’s exponentially-scaled age span to account for the time that a paper had to accumulate citations over the years $x$ and derive the Total Citation Impact (TCI) of a paper as
$\text{TCI}=\frac{\textrm{TCS}}{(\max(x)-\min(x)+1)\cdot e^{\alpha}}$
where $x$ are the years in the citation series $c$ and $\alpha$ is an exponential factor that adjusts for the fact that citation accrual follows a power law. For estimating the parameter $\alpha$, we analyze all citations ever made in the CHI proceedings. For each cited paper, we construct the citation series (citations over the years) and fit a powerlaw distribution to these series. This results in an estimate of $\alpha=3.63$. TCI is the first component of the Milestone Coefficient. The second component is the Normalized Citation Impact (NCI) which provides a measure of a paper’s citation performance relative to all other citations to papers in a proceedings years $x$. NCI is defined as the sum of this relative impact:
(2) $\text{NCI}_{j}=\sum_{x}\frac{\text{citations}_{jx}}{\sum_{i}\text{citations}_{ix}}$
We define the Milestone Coefficient (MC) of a paper as
$\text{MC}=\text{SIGN}~{}\times~{}\sqrt{\text{TCI}^{2}~{}\times~{}\text{NCI}^{2}}$
The euclidean norm defines MC as the distance to the zero point (0,0) in a coordinate system spanned by TCI and NCI. This can be explained as follows. A high-performing milestone will show exceptional performance relative to other papers (NCI), but also when adjusted for its own age (TCI). The latter accounts for natural growth in citations that every paper experiences over its lifetime. This natural accrual of citations, however, does not necessarily make a paper a milestone. A high-performing milestone will demonstrate strong performance when compared to its own growth in citations (high TCI). If either TCI or NCI is low, the overall MC is correspondingly low, emphasizing that both relative citation impact (NCI) and age-adjusted citation performance (TCI) are essential for a paper to be considered a milestone. The euclidean norm reflects a paper’s significance across both dimensions. The milestone coefficient is deliberately not scaled to fall in the unit interval [-1, 1] to acknowledge the dynamic and evolving nature of citations. Instead, a large positive milestone coefficient indicates a strongly performing and growing milestone, and a negative MC indicates a milestone in decline. We apply the milestone coefficient to the 3,000 milestones most cited at CHI and plot the results later in Figure 6.
4. Results
4.1. Milestones cited by the CHI Community (RQ1 and RQ2)
The CHI community’s 30 most cited milestones with year of publication and citation count are listed in Table 1. Half of the 30 milestones most cited at CHI are published at the CHI Conference, the other half in other conferences, journals, or books. The distribution of the CHI community’s citations to milestones is long-tailed. Very few milestones received many citations, and there is a stark difference between CHI’s top performing milestones and the rest of the milestones (cf. Table 1). This demonstrates how much the CHI community’s attention rests on a few top-performing papers. Table 1 also highlights how relatively few citations the milestones are receiving from CHI authors. For instance, the thirtieth paper in the top-30 list has received only 75 citations from the CHI community since its publication in 2006. Expanding this list to the top-300 most cited papers, the papers in 300th place each received only 34 citations from the CHI community. The following section provides a description of CHI’s top-30 most cited milestones.
4.1.1. Milestone description (RQ1)
The 30 milestones listed in Table 1 represent key contributions that have significantly influenced the landscape of HCI research. Together, these milestones reflect the diverse advances that have shaped HCI research and practice. The following section presents a typology of milestones based on their citation performance.
Figure 2. Relative number of citations to CHI’s overall top-100 most cited milestones over time. A clustering analysis reveals three types of milestones in CHI: one “super milestone” continues to grow in popularity (red), some past super milestones are fading from the CHI community’s memory (light blue), and some contemporary milestones are receiving continued attention, although at lower levels (blue).
[…]
Line plot of citations from CHI articles to the top-100 milestones over time, graphically showcasing three different types of milestones: contemporary milestones, super milestones, and fading super milestones.
Table 1. Milestones most cited in the ACM CHI Proceedings (1981–2024).
[…]
$\dagger$ Since our focus is on the CHI community, the citation count listed in this table only includes citations originating from CHI papers.
Table 2. Contributions of the 30 milestones most-cited by the CHI community.
[…]
Description and contributions of each of the 30 milestones most-cited by the CHI community. No. Year Milestone Description Contribution 1. (2006)   Braun and Clarke (Braun and Clarke, 2006) Thematic Analysis, a method for analyzing qualitative data method 2. (1988)   Hart and Staveland (Hart and Staveland, 1988) NASA Task Load Index (NASA-TLX), a tool for measuring perceived workload method/tool 3. (2011)   Wobbrock et al. (Wobbrock et al., 2011) Aligned Rank Transform (ART) for non-parametric factorial analyses through ANOVA procedures method 4. (2014)   Zimmerman and Forlizzi (Zimmerman and Forlizzi, 2014) Research through Design as a way to generate knowledge by reflecting on design practice method 5. (2010)   Bardzell (Bardzell, 2010) Feminist HCI, a critical agenda that considers gender, power, and identity in design framework/agenda 6. (2012)   Braun and Clarke (Braun and Clarke, 2012) Thematic Analysis, a method for analyzing qualitative data method/guidelines 7. (2019)   Braun and Clarke (Braun and Clarke, 2019) Reflexive Thematic Analysis, a clarification of Braun and Clarke’s initial method method/guidelines 8. (2006)   Hart (Hart, 2006) A reflection on the NASA-TLX’s two decades of impact method/tool 9. (2019)   McDonald et al. (McDonald et al., 2019) Guidelines on reliability and inter-rater reliability in qualitative research guidelines 10. (2010)   Li et al. (Li et al., 2010) A stage-based model of personal informatics systems model 11. (2003)   Hutchinson et al. (Hutchinson et al., 2003) Technology Probes, an expansion of Cultural Probes method 12. (2005)   Sengers et al. (Sengers et al., 2005) Reflective Design, an approach to question underlying assumptions in technology framework 13. (1996)   Brooke (Brooke, 1996) System Usability Scale (SUS), a tool for assessing usability method/tool 14. (2003)   Gaver et al. (Gaver et al., 2003) Ambiguity as a Resource for Design framework/concept 15. (2017)   Schlesinger et al. (Schlesinger et al., 2017) Intersectional HCI, a critical perspective advocating for designs that address the complexities of gender, race, and class framework 16. (1997)   Ishii and Ullmer (Ishii and Ullmer, 1997) Tangible Bits, a vision for seamless interfaces that blend the digital and physical worlds, laying the groundwork for tangible interaction concept 17. (2010)   DiSalvo et al. (DiSalvo et al., 2010) Sustainable HCI, urging the field to consider environmental impacts framework 18. (2010)   Irani et al. (Irani et al., 2010) Postcolonial Computing, a framework for examining design and development through postcolonial theory framework 19. (2014)   Rooksby et al. (Rooksby et al., 2014) The concept of “lived informatics”, examining how people integrate personal tracking into their daily lives model 20. (2019)   Amershi et al. (Amershi et al., 2019) Guidelines for Human-AI Interaction, offering best practices for designing user-centric AI systems guidelines 21. (2013)   Dunne and Raby (Dunne and Raby, 2013) Speculative Design, a method using fictional scenarios to explore alternative futures framework/concept 22. (1999)   Gaver et al. (Gaver et al., 1999) Cultural Probes, a method using evocative artifacts to engage people in the design process method 23. (2015)   Epstein et al. (Epstein et al., 2015) Lived Informatics Model, a development of Rooksby et al.’s method model 24. (2016)   Azmandian et al. (Azmandian et al., 2016) Haptic Retargeting, a technique for creating more immersive interactions in VR experiences method 25. (2009)   Wobbrock et al. (Wobbrock et al., 2009) User-defined gestures, underscoring the importance of user preferences in surface computing interactions method 26. (2013)   Follmer et al. (Follmer et al., 2013) inFORM, a shape display that allows for dynamic physical affordances through actuation, demonstrating the potential of tangible interfaces in HCI method 27. (1993)   Kennedy et al. (Kennedy et al., 1993) Simulator Sickness Questionnaire, a method for quantifying simulator sickness method/tool 28. (2011)   Bostock et al. (Bostock et al., 2011) D3 (Data-Driven Documents), a software library for creating interactive web visualizations tool 29. (2003)   MacKenzie and Soukoreff (MacKenzie and Soukoreff, 2003) Standardized phrase sets for evaluating text entry techniques as benchmark for assessing typing performance method/tool 30. (2006)   Charmaz (Charmaz, 2006) A practical guide to Constructing Grounded Theory method
Figure 2 tracks citations from the CHI community to its 100 most cited milestones over time, relative to all citations made in a given CHI proceedings year. Our cluster analysis of these 100 citation series finds three different types of milestones being cited at CHI:
  1. (1)
      Super milestones (1%): One milestone continues to grow in popularity, although its citations from CHI authors may be approaching saturation in recent years. In the 2024 CHI proceedings, this milestone received 0.2% of all citations made in CHI papers. Example:
  2. (2)
      Fading super milestones (4%): Four milestones were very popular in the past, with one receiving over 0.5% of all citations in one proceedings year. However, these milestone have been fading from the CHI community’s memory. Examples:
    • Card et al.’s “The Psychology of Human-Computer Interaction” (Card et al., 1983)
    • Ishii and Ullmer’s “Tangible Bits” (Ishii and Ullmer, 1997)
    • MacKenzie’s “Fitts’ law as a research and design tool” (MacKenzie, 1992)
  3. (3)
      Milestones (95%): Some milestones are receiving continued attention, although at levels much lower than prior super milestones. Examples:
Next, we turn our attention to analyzing the different contribution types the milestones make. This is crucial for understanding what type of work is being appreciated in the long run.
4.1.2. Milestone contributions (RQ2)
Table 2 provides a brief description of the unique contribution of each of the 30 milestones most cited at CHI. Our analysis identifies seven key types of milestone contributions: method, tool, framework, agenda, guidelines, model, and concept (see Table 1 and Table 3). These seven categories encapsulate the key ways in which the top-30 milestones have impacted and advanced the field of HCI.
We compare the key contribution types of the 30 milestones with Wobbrock and Kientz’s seven types of research contributions in HCI (Wobbrock and Kientz, 2016) and the Contribution Types for CHI 2024 (ACM SIGCHI, 2024) (see Table 3). The most prominent contribution shared with existing typologies is the method contribution, which directly corresponds to Wobbrock and Kientz’s methodological contribution. More than half of the top-30 milestones make methodological contributions (see Table 1).
However, many other contribution types in HCI are not represented in the CHI community’s milestone contributions. For instance, prototypes and systems are a common type of contribution in HCI research, with several hundred papers designing, implementing, and evaluating systems and prototypes each year. However, this type of contribution is not represented in the list of key milestone contributions. Also not represented are less common contribution types, such as literature survey, dataset, and replication.
Table 3. Comparison of milestone contribution types with existing contribution types in HCI.
[…]
Comparison of milestone contribution types with existing contribution types in HCI. Contribution Description with examples Wobbrock and Kientz ((2016))* ACM SIGCHI ((2024)$\dagger$ Method A systematic approach or technique used for conducting research or analysis in HCI.
Examples: qualitative methods (e.g., thematic analysis), quantitative methods (e.g., aligned rank transform)
Methodological Methodology
Tool A practical instrument or software designed to perform specific tasks or measurements in HCI studies.
Examples: survey tools (e.g., NASA-TLX, SUS), software libraries (e.g., D3)
Artifact Artifacts
Framework A comprehensive structure outlining key concepts in HCI which organizes knowledge and provides a structured approach for HCI research.
Examples: Feminist HCI, Reflective Design
Theoretical Theory
Agenda A call for future research or practice directions in HCI.
Example: Feminist HCI agenda
Opinion Argument
Guidelines Prescriptive recommendations or best practices for designing or evaluating human-computer interaction.
Example: Guidelines for inter-rater reliability
Methodological Methodology
Model An abstract representation detailing the components and processes of a system or phenomenon in HCI.
Example: Model of personal informatics
Theoretical Theory
Concept A foundational idea or theory that introduces new perspectives or approaches to HCI design and research.
Example: Ambiguity
Theoretical Theory
* Research contribution types in HCI (Wobbrock and Kientz, 2016): Empirical, Artifact, Methodological, Theoretical, Dataset, Survey, Opinion $\dagger$ Contributions to CHI (ACM SIGCHI, 2024): 1) Development or Refinement of Interface Artifacts or Techniques, 2) Understanding Users, 3) Systems, Tools, Architectures, and Infrastructure, 4) Methodology, 5) Theory, 6) Innovation, Creativity, and Vision, 7) Argument, 8) Validation and Replication
4.1.3. Case study: The rise of CHI’s super milestone
Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006) is a unique “super” milestone (cf. Figure 5 and Figure 6). What makes this milestone stand out is not only its absolute citation count, but also its speed of adoption in the CHI community. Braun and Clarke’s method for qualitative analysis of data was first published in the year (2006) and was immediately useful to researchers in HCI, resulting in rapid adoption in the CHI community (cf. Figure 2 and (Bowman et al., 2023)). In fact, Thematic Analysis occupies not one but three of CHI’s 30 most cited milestones (see Table 1). Together, these milestone papers have amassed 1024 citations from CHI papers. While this number pales in comparison to the total number of citations that these milestone papers have received from other sources, it does demonstrate the popularity of Thematic Analysis in the CHI community. Braun and Clarke’s milestones have received almost four times more citations than the second most cited milestone at CHI (see Table 1).
In the year 2024, Braun and Clarke’s most-cited milestone (Braun and Clarke, 2006) accounts for about 1 in 500 citations (0.2%) in the CHI proceedings. Given that CHI papers include an average of about 88 references in 2024 (Oppenlaender, 2024), this means we can estimate that, on average, about one in six (5.7) CHI papers cite Braun and Clarke’s work. At present, no other milestone demonstrates such an exceptional performance. In the following section, we examine how the CHI community’s attention to this and other milestones has evolved over time.
4.2. How has the CHI Community’s Attention to its Milestones evolved over time? (RQ3)
Forgetting curves graphically depict the attention of the CHI community to past papers and milestones (see Figure 1). A visual inspection of the forgetting curves yields the insight that the forgetting curves have become “smoother” over time. Each year, there are fewer visible peaks in the forgetting curves and, thus, fewer milestones are being cited in the ACM CHI proceedings. In other terms, papers published today are less likely to cite older papers.
Figure 3. CHI’s forgetting curves have become “smoother” over time, lacking the visible peaks that are characteristic of milestones.
[…]
Line plot of variance of derivative and total variation of the CHI forgetting curves.
We can quantify this visual effect by calculating and plotting the variance of the derivative and the total variation of each of the forgetting curves (see Figure 3). The variance of the derivative measures the variability of the rate of change (derivative) in the forgetting curves. The total variation represents the cumulative amount of change over the forgetting curve. The plot in Figure 3 shows that the forgetting curves have become “smoother” over time, providing graphical evidence of a fading memory of milestones in the CHI community.
Figure 4. Cumulative citations to the prior 300 most cited milestones from selected CHI proceedings years.
[…]
Cumulative citations to the prior 300 milestones from selected CHI proceedings years.
In the 2024 Proceedings of CHI, a significant portion of the community’s attention is directed toward more recent contributions. To understand this trend, we calculated the cumulative percentage of citations from selected CHI proceedings to the 300 most cited milestones published in years prior (see Figure 4). The results highlight the concentration of citations to recent milestones. Notably, 15.4% of the citations point to works published in 2023. This cumulative percentage increases to 27.5% when including works from 2022 and 2023, and further rises to 60.1% when considering all works from 2018 onwards. By 2010, the cumulative percentage reaches 79.6%, and by 2005, it climbs to 89.9%. The bulk of the CHI community’s attention rests on papers published in a time window of between 5 and 15 years. This indicates a strong emphasis on citing more recent milestones within the CHI community.
Figure 5. Absolute (top) and relative (bottom) number of citations from selected CHI proceedings to the 300 most cited milestones published in years prior to the given proceedings year. Note that for 2002 and 2004, the number of milestones is marginally lower. Standard deviation, kurtosis, and skewness values are listed in each plot. Looking at the panels, we can identify that while the absolute number of citations to prior milestones has increased, the relative number of citations to prior milestones has decreased significantly. In other terms, milestones have lost influence over time at CHI, relative to all other citations in a given proceedings year.
[…]
A plot depicting how the CHI community’s citations to the 300 most cited prior milestones have developed over time. Looking at the panels, we can note that milestones have lost influence over time at CHI, relative to all other citations in a given proceedings year.
The picture is more nuanced when considering the evolution of citations at CHI (see Figure 5). In Figure 5, we can observe that the CHI community’s citations have always been skewed (see Figure 5), favoring more recent milestones over older ones. An exception to this is the CHI community’s preferential attachment to its super milestone which is clearly visible in recent proceeding years. But while there have been small differences over the years, there is no clear trend toward citation distributions becoming more skewed in recent years.
There is, however, a notable shift between the absolute and relative number of citations to milestones in recent years. In absolute terms, citations to the top-300 of CHI’s milestones have increased, including citations to older milestones. The growth in the absolute number of citations in Figure 5 can be attributed to the extreme growth of references in CHI papers since 2016, as observed in prior work (Oppenlaender, 2024). In the 2024 proceedings alone, the CHI community made 3,951 citations to the prior top-300 milestones, compared to only 359 citations ten years earlier in the 2014 proceedings. In relative terms, the 2024 number of citations to milestones pales in comparison to the total number of citations made in that proceedings year (93,639 citations). The top-300 prior milestones only account for a small fraction of citations in 2024 (4.2%). This demonstrates that while the absolute number of citations to milestones has increased, the relative number of citations to milestones has declined. For comparison, the milestones published in the prior five years each received about 3% of all citations in 2002 (top left in Figure 5). In 2024, the relative number of citations to prior milestones is almost ten times lower (bottom right in Figure 5). This highlights that in relative terms, older milestones are being cited less in the CHI community.
Figure 6. Milestone coefficient (MC) for the 3,000 papers most cited by CHI authors. The year corresponds to the milestone’s publication year, and the size of the bubble reflects the total number of citations to the milestone from CHI papers (over all CHI proceedings). Papers below the horizontal line tend to being forgotten by the CHI community over time, while papers above the line continue to grow in citations. The distance from the dashed horizontal line indicates the milestone’s impact, with high-performing milestones having a greater positive distance from the line. The plot also demonstrates the exceptional performance of the super milestone published in 2006.
[…]
Milestone co-efficient plotted for the 3,000 papers most cited at CHI.
We apply the milestone coefficient (MC) to the top-30 milestones (as listed in Table 1). The MC in Table 1 is negative only for one milestone (D3 (Bostock et al., 2011)) which has seen a decline in citations from the CHI community in recent years. All other milestones in the top-30 list are still trending toward growth in citations. To extend this analysis and derive insights on the CHI as a whole, we apply the milestone coefficient to the 3,000 papers most cited by CHI authors. For these papers, the milestone coefficient ranges from -0.0102 to 0.0768. Figure 6 plots the MC for the top 3,000 most cited papers. Each bubble represents one paper in its year of publication. The size of the bubble represents the total number of citations to this milestone. The exceptional super milestone is clearly visible in the year 2006, achieving an MC of 0.0768 (the highest MC in the dataset). The second most cited milestone, the NASA TLX (Hart and Staveland, 1988), is visible in 1988. From Figure 6, it can be derived that the number of milestones has grown over the years. This is unsurprising, given the growth of the CHI proceedings and the growth in citations (Oppenlaender, 2024). However, the figure also demonstrates that many recent milestones struggle to receive the same attention from the CHI community as the past super milestones. Some of the recent milestones even trend towards a decline in citations, and the number of milestones that are being ‘forgotten’ has increased in the last two decades.
Our analysis, so far, has been focused on citations and papers as unit of analysis. In the following section, we shift the focus to the milestone authors.
4.3. Does the CHI Community show a Preference for its Milestone Authors? (RQ4)
The CHI community is relatively large, with over a thousand authors attending the CHI Conference. While attended by many first-timers, some authors at CHI are repeat authors (Kaye, 2009). The Matthew effect in science (Merton, 1968) describes the tendency of successful authors to be cited more often. This accumulated advantage may also be present in the milestones at CHI, and we expect many milestones authors have authored more than one milestone.
Figure 7. A plot depicting the number of authors who have authored one or more papers among the CHI community’s 300 most cited milestones. This includes articles published at CHI and external venues. While the majority of authors have only written one of the top-300 milestones, about 20% of the authors have (co-)authored multiple of the top-300 milestones. A logarithmic scale is used in this plot to make small values visible.
[…]
A bar chart depicting the number papers that authors have (co-)authored among the top-300 most cited milestones at CHI. While the majority of authors have only written one of the top-300 milestones, about 20% of the authors have (co-)authored multiple of the top-300 milestones.
To investigate this, we identify the authors of the 300 milestones most cited in the CHI community (over all proceedings years) and count the number (see Figure 7) and cumulative percentage (see Figure 8) of milestones these authors have (co-)authored. We find that while the majority ($n=615$, 80.3%) of the milestone authors have only authored one of the top-300 milestones, 151 authors (19.7% of the top-300 milestone authors) have (co-)authored more than one of the top-300 milestones. The 50 authors who have (co-)authored three or more of CHI’s 300 most cited milestones are listed in Figure 8. Chief among them are Phoebe Sengers, Paul Dourish, and Jacob O. Wobbrock who each have (co-)authored eight of the CHI community’s 300 most cited milestones. This is followed by John Zimmerman (7 milestones), Jodi L. Forlizzi (6 milestones), and others (see Figure 8). Together, the 50 authors have (co-)authored about 20% of the 300 milestones most cited by the CHI community (see secondary axis in Figure 8).
Figure 8. Fifty authors have published three or more of the 300 milestones most cited in the CHI community (teal bars and lower x-axis). The 50 authors have (co-)authored about 20% of the 300 milestones (red line and top x-axis). The milestones publication years are listed in parentheses.
[…]
Authors having authored papers three or more milestones in the 300 papers most-cited by the CHI community. The 50 authors have (co-)authored about 20% of the 300 milestones (red line and top x-axis).
5. Discussion
The CHI community continues to grow in size and impact. In this paper, we presented meta-research on how the community operates as a scientific field. The way how we cite our milestones is not only an intellectually satisfying exercise, it also concretely explains what kind of work goes appreciated at large in the community.
5.1. The CHI Community’s Milestones
What makes a milestone? And what makes a paper stand the test of time? Our research gives empirical indications of what is valued at CHI, as expressed through citations to milestone papers. In our analysis, we found that milestones cited at CHI, and – as a proxy, in HCI – make mainly methodological contributions. These contributions deeply engage with the field’s history and innovate over the state-of-the-art, resulting in a long-lasting impact in HCI.
This contrasts with some of the contribution types in HCI that are not represented in the list of top-cited milestones. In particular, milestones are not being cited for contributing prototypes and systems, which is one very popular contribution type at CHI. Spurred by the availability of large language models, it is now easier than ever to implement prototypes in HCI and write a paper about them. We can observe that some authors rush from one quickly implemented HCI study to the next (Noorden and Perkel, 2023). The methodological acceleration in HCI (Thomas et al., 2024) shifts our attention away from things that may have a long-lasting impact in HCI in favor of quick results. Some researchers have raised concerns about the quick-win approach found in some HCI research, calling it concerning and detrimental to the field of HCI (Arawjo, 2024). HCI has a strong case of “shiny object syndrome”, where short-lasting projects are prioritized over time-persisting milestones. We believe it is not likely that the increasingly popular “thin-wrapper” implementations with language models will become part of the set of milestone contributions in HCI. It is important for us to not forget the past milestones and their contributions, as they inform our methodologies and values in HCI.
Milestones also do not include some other HCI research contribution types, including literature reviews, datasets, and opinions. In particular, we note that literature reviews are a very young contribution type at CHI. Oppenlaender found that literature reviews only became a topic at the CHI Conference after page restrictions were lifted in 2016, but have since grown in popularity (Oppenlaender, 2024). While it has also become more common to cite datasets and code repositories in CHI papers (Oppenlaender, 2024), datasets are not among the top milestones cited at CHI. Opinions are, perhaps, the least known contribution type and, for this reason, not common at CHI.
Another type of contribution type that is not represented in the milestones are replications. Much of the research in HCI is concerned with gathering primary data, which raises concerns for the replicability of HCI studies. This “replication crisis” in science is a well-known problem (Baker, 2016; Cockburn et al., 2020; Ioannidis, 2005; Korbmacher et al., 2023). In HCI, several authors have studied replicability (Hornbæk et al., 2014; Echtler and Häußler, 2018), but replication studies have not made it among CHI’s top-cited milestones. As Wilson et al. noted in their RepliCHI workshop summary at CHI 2014, HCI researchers “have almost no drive and barely any reason to consider investigating the work of other HCI researchers” (Wilson et al., 2014).
This preference for certain contribution types at CHI has implications for authors. Scientists who want to contribute to HCI’s long-lasting intellectual heritage should consider developing novel methods and frameworks that advance the field. However, we acknowledge this is easier said than done, because these novel methods compete with existing methods, and novel methods may not receive attention from the community. One example is Scollon and Scollon’s Nexus Analysis (Scollon and Scollon, 2004). This research method has first been presented in (2004), but only a limited number of studies at CHI use this method (e.g., (Iivari, 2019) and (Norouzi et al., 2021)).
These examples reflect how the CHI community prioritizes and balances its attention to past and present milestones. Our work addresses a critical issue in the CHI community: the balance between innovation and the retention of foundational knowledge. Older milestones may, over time, get pushed into the long-tail of the citation distribution where the milestones do not receive enough attention to rise to super milestone status. This also suggests a shifting focus in the CHI community away from foundational works in favor of more recent contributions. However, due to the extreme growth in the number of citations made in CHI papers, newer papers struggle to receive the collective attention of prior super milestones, and it is unclear whether newer papers will be able to rise to the level of fame of older super milestones. It seems that most recent milestones, while relevant to some, fail to capture the CHI community’s collective attention compared to the success of past milestones.
The “sweet spot” of relevance for papers cited at CHI is 5–15 years back in time. The 5–15 year citation range likely reflects a balance between novelty, relevance, accessibility, and consolidation of knowledge. Research within this range has typically undergone sufficient validation and follow-up, making it both trustworthy and broadly applicable, while remaining relevant to current advances. This time frame allows foundational ideas to influence new work, often through citation cascades that amplify visibility and integration into broader academic and practical contexts. Ideas from this period are also more likely to have been incorporated into curricula and practice, making them accessible and influential to a wide audience. Similar patterns are seen in other disciplines, such as physics and medicine, where impactful research takes years to be validated, cited, and adopted into subsequent innovations. Conversely, older work may become less cited as it is superseded by newer paradigms or rendered less accessible. This highlights the dynamic interplay between innovation, validation, and knowledge dissemination in shaping research impact.
Another trend defining the dynamics of milestones is the growth in citations and publications at CHI. As the CHI Conference evolved and scaled in diversity and size, milestone papers that concern the whole CHI community are less frequently occurring. This development creates a dilemma for the CHI Conference: the conference has, arguably, become too large and diverse to have common denominators and shared research interests. What matters to one sub-community in CHI might be uninteresting to other sub-communities.
A community that does not celebrate its past milestones is a community that speedruns from one ephemeral project to the next. If milestones are no longer valued and cited, and if the field fails to discover milestone papers in its publications, then, due to the growth of publications and the shifting attention of the field, it is becoming less and less likely every year for CHI to birth new milestones. This is arguably concerning and endemic of the fractured nature of the CHI community in many sub-communities around different highly diverse topics (Liu et al., 2014; Oppenlaender and Hämäläinen, 2023).
These trends are indicative of a community at a crossroads between remembrance and forgetfulness, a community that while honoring and valuing past contributions, is gravitating towards more recent work. Whether this is a problem or natural evolution is, of course, subjective. These issues are known and have, of course, also been grappled with earlier, e.g. through the CHI2030 Vision Task Force (2018–2019) that sought to address the future of the CHI Conference.
5.2. Matthew Effect and Authoring at CHI
We demonstrated how certain prolific authors contribute a lot to the community milestones. The 50 most highly-cited authors have (co-)authored about 20% of the top-300 milestones cited at CHI, as depicted in Figure 8. Reasons for this concentration of milestone authorship may vary. Obviously, the work itself by these authors is of high-quality and constitutes advances worth remembering in new and upcoming research. And some preferential attachment is normal in a research community. On the other hand, some of the citation patterns might be explained through other forms of influence: large researcher networks, access to labs with plenty of funding and, thus, future researchers citing the past work from the same lab, etc. These are all factors that we should, in all fairness, recognize. It may be rather difficult to attain a milestone status for one’s work, if it originates from an underdog position and from a relatively unknown laboratory. In any case, this concentration of milestone authorship suggests an influential core within the CHI community.
What does this mean for authors at CHI? The cumulative advantage of existing milestones makes it difficult for emerging authors to publish new milestones. While it may arguably be easier for authors to contribute an impactful milestone in one of CHI’s many sub-communities, authoring a milestone that concerns the whole community is hard, as new milestones compete with the CHI community’s attention and established milestones. CHI’s attachment to its one super milestone is the best example for this accumulated advantage. The accumulated advantage and focused attention of the community on some highly performing milestones, we argue, creates barriers to entry for emerging milestones and their authors. One could speculate on whether the CHI community is saturated with milestones, which would further contribute to barriers of entry. Braun and Clarke’s work, for instance, seems to receive fewer citations (relative to all citations in the CHI Proceedings) in recent years, which could indicate that the CHI community has reached saturation with this important research method. This suggests that the CHI community may be at a point where established milestones overshadow the emergence of new, impactful contributions. In the following case study on CHI’s super milestone, we explore how this influential milestone navigated these challenges, setting a precedent for how future milestones might overcome these barriers.
5.3. Case Study on CHI’s Super Milestone
Our case study on CHI’s super milestone invites reflection on what makes a milestone enduring—beyond just time or relevance to the field. What criteria elevate a contribution to super milestone status and what does the CHI community implicitly value? It is clear that Thematic Analysis provided a versatile method for qualitative data analysis. This methodological guidance and research method proved to be a highly useful tool in the HCI researcher’s toolbox.
However, Braun and Clarke have more recently acknowledged that many researchers apply their method not as it was originally intended (Braun et al., 2019; Braun and Clarke, 2019). The authors themselves now advocate for the ‘fading’ of their own milestone in favor of their more recent works on Reflexive Thematic Analysis that explain how their original idea was intended. In many cases, researchers only loosely apply the method as an umbrella for their (unstructured) qualitative investigations or they apply different versions of the approach, as delineated by Braun et al. (Braun et al., 2019). Kaltenhauser et al. found the same issue in their review of autoethnographies, noting that “existing methods are often executed and interpreted with much variety” (Kaltenhauser et al., 2024). This speaks to how researchers often (mis-)appropriate methods and milestones.
Bowman et al. presented a scoping review on the use of Thematic Analysis in Healthcare HCI, noting that the “departures from what is advocated as quality [Thematic Analysis] practice” constitute “a change in research practices” (Bowman et al., 2023). In other words, the appropriation of the method by HCI researchers is testament to the innovation and continued evolving nature of the Thematic Analysis method. Like a spoken language that changes and evolves over time, a research method will be adapted and evolved by people who apply it. But how can a research community refer to this evolved version of the method, if the method deviates significantly from its original version? Clearly, many authors still attribute their evolved use of the method to the original milestone. This is where articles such as Braun et al.’s later work (Braun et al., 2019) come in, because this article takes review of how the Thematic Analysis milestone was applied, and—crucially—still allows researchers to continue their use of the appropriated method, should they choose to. For instance, one can still use a “codebook” approach to Thematic Analysis, and the correct milestone to cite for this approach is (Braun et al., 2019).
5.4. The Milestone Coefficient: A New Lens on Impact
We introduced the Milestone Coefficient as a novel approach to quantifying the significance of contributions on a continuous scale. The Milestone Coefficient is intended as a tool to identify milestones and reassess what constitutes a milestone in a research community, offering a fresh perspective that goes beyond traditional citation counts. More specifically, the Milestone Coefficient considers two factors. First, it considers the relative impact of a milestone compared with other papers in the field. Second, it considers the performance of a milestone against its own growth in citations. Scholarly papers will naturally accumulate citations over their lifetime, but this does not necessarily make them milestones. A milestone will show exceptional performance against its age-adjusted citations.
Many research indicators, such as the h-index, focus on citations. Citations as a metric for evaluating papers are problematic for several reasons. One reason is that milestones are often cited for the wrong reasons, as discussed in the previous section. Consequently, many milestone papers accumulate citations only for being a milestone. Another reason is that citations have become extremely important to a researcher’s career (Edwards and Roy, 2017). A recent survey of over 30,000 faculty members of the 10 highest ranked universities globally found that Google Scholar is the most popular source of information to evaluate a researcher’s career (Ibrahim et al., 2024). The same study demonstrated that citation counts in Google Scholar can be manipulated, which has also been demonstrated by other studies (e.g., (Labbé, 2010)). Citations, therefore, are a shallow proxy for research success and can be manipulated.
The Milestone Coefficient is by no means intended as a tool for research assessment. The MC, as yet another bibliometric metric, will not help with addressing the current problems with bibliometrics and academia. Further, only very few authors will—by definition—become milestone authors. For this reason, the Milestone Coefficient would be ineffective for the purpose of evaluating the career progress of most researchers. Rather, our intention with the Milestone Coefficient is to provide a tool and quantitative means for identifying milestones, discussing, and comparing the performance of milestones, going beyond a discrete typology of milestones. We hope this coefficient will provoke a discussion on the works that have had the greatest impact at CHI and about what we value in the HCI research community.
5.5. Limitations and Future Work
We acknowledge limitations to the validity of our research. First, while the ACM CHI Conference is a leading venue in HCI research and, thus, could be viewed as a suitable proxy for investigations on the wider field of HCI, our investigation is limited to our subject of study: The CHI community. It is our intention to understand the citation behavior within this particular community, and how the community has evolved over time. Future work could expand to other research communities, providing valuable comparison points.
Second, our work examines how the CHI community cites its milestones. It is, however, possible that the CHI community favors derivative works over milestones. Metaphorically, authors of derivative works stand on the shoulders of giants, but by becoming giants themselves, they may crush what is underneath them. One example of this is Fitts’ Law (Fitts, 1954). Fitts’ highly influential work was published in (1954) and inspired much research in HCI. However, Fitts’ success was shared by other authors who presented studies on Fitts Law. While we do not find strong evidence for other instances of this in our research, future work could investigate how newer milestones and derivative works displace attention to older milestones.
Third, we acknowledge that our method of identifying milestone papers and authors are imperfect. We uniquely identify papers by their title and publication year. We use regular expressions to extract the paper titles. While we use normalization (such as lower casing and removal of punctuation and white spaces), our simple string matching approach is limited in uniquely identifying papers. References in CHI papers are noisy and may contain human mistakes, such as errors and typos. Our method cannot address these human errors. In a similar vein, author name disambiguation is a common problem in bibliometric analyses that is impossible to solve completely (Bartneck and Hu, 2009; Kaye, 2009). For instance, authors may use different versions of their name, or they may legally change their name entirely. While we did our best to develop a set of regular expressions to capture different notations of author names, human errors can not be fully accounted for.
Fourth, the growth in the number of papers and the increase in the number of references included in CHI papers may contribute to gradual smoothing of the forgetting curves over time. This would not be an issue if changes were proportional (such as a proportional increase in the number of references and papers). However, as described by Oppenlaender, there was a shift in citation behavior in 2016 that constitutes a non-linearity in the CHI community’s trajectory (Oppenlaender, 2024). The forgetting curves should be interpreted with this non-linearity in mind.
Last, our methodology focuses on total citations over the entire ACM CHI conference. This methodology is not likely to identify milestones that have been popular for a very short period of time but since have faded completely from CHI’s collective memory. This would not allow the milestones to have accrued enough citations to be noticable in our methodology. We leave this and investigations of milestone impact in CHI’s sub-communities to future work.
6. Conclusion
We presented an analysis of how the CHI community is, over time, shifting its attention away from its past milestone papers. Forgetting curves provide graphical evidence that the relative attention of the CHI community to its milestone publications is decreasing over time. The attention span of the CHI community narrowly rests on milestone papers published in about ten year window. While this is a testament for the dynamism and renewal of research at CHI and in the field of HCI, it raises concerns about the value, sustainability, and meaningfulness of CHI research. The HCI field’s key contributions consist of methodological and conceptual contributions that move the field forward. However, it is unclear whether recent milestones can achieve the fame of past milestones in HCI.
Our findings and the Milestone Coefficient offer practical insights for researchers and policymakers in HCI. We hope our work will provide insights into the dynamics of research recognition and longevity in HCI, provoke a critical reflection in the CHI community on its fundamental values, and foster discussion on sustainable strategies forward.